The SAT and ACT college entrance exams are taken by thousands of students each year. The mathematics portions of each of these exams produce scores that are approximately normally distributed. In recent years, SAT mathematics exam scores have averaged 480 with standard deviation 100. The average and standard deviation for ACT mathematics scores are 18 and 6, respectively.

(a) An engineering school sets 560 as the minimum SAT math score for new students. What percentage of students will score below 560 in a typical year? (Round your answer to two decimal places.)

___________%

(b) What score should the engineering school set as a comparable standard on the ACT math test?
Answer #1

a)
P(X <= 560) = P(z <= (560 - 480)/100)
= P(z <= 0.8)
= 0.7881

78.81 %

b)
ACT score,
x = mean + z*sigma
x = 18 + 0.8*6
x = 22.8